There are following options available on the market: a. CALL with a strike price of 100 PLN and premium of 5 PLN b. PUT with a strike price of 100 PLN and premium of 10 PLN Is an arbitrage possible explain your strategy? Would it be possible if you could either buy or write above options with the same characteristic- explain your strategy
Answer #1

Using put call parity
S+P=C+K/(1+r)^t
Here, S+P=110
K/(1+r)^t<100
as K=100 and r>0 and t>0
So, S+P>C+K/(1+r)^t

Hence, we sell overvalued side
in this case lhs
and buy undervalued side
in this case rhs

Yes arbitrage is possible

Sell stock
Sell put
Buy call
Buy a bond with face value equal to 100